# Human vs Machine: Comparison of Student-Generated and AI-Generated Educational Content

### v1

@@ Abstract
Keywords Learnersourcing  $\cdot$ Large language models  $\cdot$ RiPPLE  $\cdot$ Codex  Generative AI  $\cdot$ Student-generated resources
@@ 3 Methods
In this section we describe how artefacts created by students participating in a learnersourcing activity were compared to those generated by an LLM. To begin, we provide an overview of the course context and the learnersourcing activity that was integrated into the course. Next, we introduce the research tool, RiPPLE, which was used to conduct the study. Lastly, we outline the study’s design, detail the data collected, and discuss the measures used for analysis.
@@ 3.2 Learnersourcing task
3.3 Research tool:RiPPLE
For our study, we made use of RiPPLE [33], which is an adaptive educational system that develops learning content through learnersourcing. This system offers three types of learning activities: create, evaluate, and practice. We now briefly describe each of these activities.
Create. Students in a course develop study resources like MCQs and worked examples. To conduct our study, we introduced a new resource type, “code example”, as previously described, to the system. To incorporate our new resource type, “code example,” we prompted students to select relevant concepts from a list of course topics when creating a new example. They were also asked to indicate the example’s level of difficulty (easy, medium, or hard) and provide a brief title. Figure 2 depicts the interface presented to students for creating a “code example” in RiPPLE.
Figure 2: The user interface of RiPPLE showing the creation process for a new “code example” artefact.
Evaluate. The study resources created by students undergo peer review by multiple peers. Reviewers use a set of criteria to rate the quality of the resource and their confidence in their rating. Additionally, they provide written feedback that justifies their rating and provides constructive feedback on how to improve the resource. Figure 3 displays the interface and evaluation criteria for assessing the quality of a code example. Using the submitted peer reviews and algorithms discussed in [37], RiPPLE makes a determination on the quality of the resource. Approved resources are added to a course repository that is accessible to all students while ineffective resources are returned to the author with feedback for resubmission. Instructors can monitor the process with the help of an AI-based spot-checking algorithm [25].
@@ 3.4 Experimental Design
To address our research questions, we conducted a study using RiPPLE in which students created code examples as part of one activity in a weekly lab session. Students were provided with six exemplars created by the teaching team to help them craft their own code examples. The code for all six exemplars was provided in the form of a single function definition. Table 1 shows all six exemplars (code is not included for space reasons). Three of these six examples (i.e. PrintAverageRainfall, SumBetween and PrintSummary) were published explicitly as part of the description of the lab activity, and the other three were made available as part of a lecture that accompanied the release of the lab resources.
@@
Figure 4 illustrates two of the new code examples that were generated by the Codex model. All 100 code examples generated by Codex were then inserted into RiPPLE and appeared in the same pool alongside the student-generated resources.
@@
Subsequently, as part of the lab activity, all of the code examples that were produced by students and generated by Codex were peer-reviewed by students using RiPPLE’s evaluation functionality. Notably, the source of the resources was not shown to students during this review phase, thus the peer evaluation was performed blind. Table 2 displays the count of the number of resources that were produced by students and by Codex, along with the total number of peer reviews submitted for these resources.

### v2

# Can We Trust AI-Generated Educational Content? Comparative Analysis of Human and AI-Generated Learning Resources

@@ Abstract
Keywords Learnersourcing  $\cdot$ Large language models  $\cdot$ Codex  $\cdot$ Generative AI  $\cdot$ Student-generated resources
@@ 3 Methods
In this section we describe how artefacts created by students participating in a learnersourcing activity were compared to those generated by an LLM. To begin, we provide an overview of the course context and the learnersourcing activity that was integrated into the course. Next, we introduce the research tool, {AnonymizedTool}, which was used to conduct the study. Lastly, we outline the study’s design, detail the data collected, and discuss the measures used for analysis.
@@ 3.2 Learnersourcing task
3.3 Research tool:{AnonymizedTool}
For our study, we made use of {AnonymizedTool} [33], which is an adaptive educational system that develops learning content through learnersourcing. This system offers three types of learning activities: create, evaluate, and practice. We now briefly describe each of these activities.
Create. Students in a course develop study resources like MCQs and worked examples. To conduct our study, we introduced a new resource type, “code example”, as previously described, to the system. To incorporate our new resource type, “code example,” we prompted students to select relevant concepts from a list of course topics when creating a new example. They were also asked to indicate the example’s level of difficulty (easy, medium, or hard) and provide a brief title. Figure 2 depicts the interface presented to students for creating a “code example” in {AnonymizedTool}.
Figure 2: The user interface of {AnonymizedTool} showing the creation process for a new “code example” artefact.
Evaluate. The study resources created by students undergo peer review by multiple peers. Reviewers use a set of criteria to rate the quality of the resource and their confidence in their rating. Additionally, they provide written feedback that justifies their rating and provides constructive feedback on how to improve the resource. Figure 3 displays the interface and evaluation criteria for assessing the quality of a code example. Using the submitted peer reviews and algorithms discussed in [37], {AnonymizedTool} makes a determination on the quality of the resource. Approved resources are added to a course repository that is accessible to all students while ineffective resources are returned to the author with feedback for resubmission. Instructors can monitor the process with the help of an AI-based spot-checking algorithm [25].
@@ 3.4 Experimental Design
To address our research questions, we conducted a study using {AnonymizedTool} in which students created code examples as part of one activity in a weekly lab session. Students were provided with six exemplars created by the teaching team to help them craft their own code examples. The code for all six exemplars was provided in the form of a single function definition. Table 1 shows all six exemplars (code is not included for space reasons). Three of these six examples (i.e. PrintAverageRainfall, SumBetween and PrintSummary) were published explicitly as part of the description of the lab activity, and the other three were made available as part of a lecture that accompanied the release of the lab resources.
@@
Figure 4 illustrates two of the new code examples that were generated by the Codex model. All 100 code examples generated by Codex were then inserted into {AnonymizedTool} and appeared in the same pool alongside the student-generated resources.
@@
Subsequently, as part of the lab activity, all of the code examples that were produced by students and generated by Codex were peer-reviewed by students using {AnonymizedTool}’s evaluation functionality. Notably, the source of the resources was not shown to students during this review phase, thus the peer evaluation was performed blind. Table 2 displays the count of the number of resources that were produced by students and by Codex, along with the total number of peer reviews submitted for these resources.
@@ 6 Conclusion
Statements on open data, ethics and conflict of interest
  • Open data: All data that was analysed for this work will be made publicly available, along with corresponding Jupyter notebook scripts.
  • Ethics statement: This research was conducted with approval of the {Anonymized - details present on title page}
  • Conflict of interest statement: There are no conflicts of interest to declare.